# MORPH: PDE Foundation Models with Arbitrary Data Modality

### v3

@@ 4 Experiments
We select datasets from three publicly available PDE benchmarks, i.e., PDEBench (takamoto2022pdebench), PDEgym (herde2024poseidon) and The Well (ohana2024well). They cover a wide variety of PDEs; the latter, however, focuses on more complex multiphysics problems. Because MORPH accommodates data heterogeneity, we demonstrate versatility in our dataset selection for both pretraining and fine-tuning. The pretraining and finetuning datasets and their shapes are listed in Table 1 and 2. Compared to existing PDE foundation models, our pretraining and fine-tuning sets are larger and more diverse. We defer more dataset-specific details to Appendix A.
@@
Pretraining and finetuning settings.
Our model accommodates both fixed–order autoregressions, $AR(p)$, and time-varying orders, $AR(p_{1:T})$, where $p$ (or $p_{t}$) denotes the autoregressive order. In this paper, we train MORPH in a self-supervised $AR(1)$ setting where we learn a map $F_{\theta}:\mathcal{X}\to\mathcal{X}$ such that for all $t\geq 0$, $X_{t+1}\approx F_{\theta}(X_{t})$ by minimizing the mean-squared error. High–order fixed AR models (e.g., $AR(16)$ in mccabe2024multiple and $AR(10)$ in hao2024dpot) often yield stronger long-horizon rollouts by reducing exposure bias via a longer receptive field. However, this design departs from the initial-value-problem (IVP) semantics that PDE foundation models aim to respect. For these reasons we adopt a one-step ($AR(1)$) formulation aligned with IVP semantics that preserves the compositional structure fundamental to PDE time evolution. Nevertheless, MORPH remains flexible and supports multi-order (fixed or varying) autoregression to improve rollout quality. For brevity, full pretraining and finetuning settings are documented in Appendix B.2, B.3.
4.1 Baselines
Models.
We compare MORPH against classical neural operators and recent PDE foundation models. As popular baselines, we use Fourier Neural Operator (FNO) (li2020fourier), a parameter-efficient spectral neural operator widely adopted for fluid surrogates, and UNet (ronneberger2015u; rahman2022u), a convolutional encoder–decoder with multiscale skip connections. Among foundation models, we include MPP (mccabe2024multiple), which demonstrates gains from multi-physics pretraining; DPOT (hao2024dpot), a denoising operator transformer with Fourier attention; and Poseidon (herde2024poseidon), a hierarchical multiscale operator transformer with shifted-window attention and lead-time-conditioned normalization for continuous-time querying.
Evaluation Metrics.
We follow the conventions of PDEBench (PB) and The Well (TW) when reporting errors on spatiotemporal PDE datasets. The Normalized Root Mean Squared Error (NRMSE) is the RMSE normalized by the $\ell_{2}$ norm of the ground truth, making it a scale-independent (dimensionless) measure and the default metric in PDEBench (takamoto2022pdebench). In contrast, the Variance-Scaled RMSE (VRMSE) normalizes the RMSE by the target’s per-snapshot standard deviation (i.e., the square root of its variance), thereby comparing errors to the intrinsic variability of the field. The Well adopts VRMSE as a primary metric (ohana2024well).
Comparison Caveats.
4.2 Main results
Our primary experiments are organized in Tables 3,4,5. In these tables, we compare baseline and SOTA models using NRMSE and VRMSE as evaluation metrics.
Datasets Models FNO UNet MPP-$S$ DPOT-$S$ MORPH-$S$
PB/TW PB/TW Ti (7M) S (30M) Ti (7M) S (30M) Ti (7M) S (30M)
Pretraining sets
1D-CFD (N) 0.095 0.36 - - - - 0.06102 0.05719
2D-DR (N) 0.12 0.84 0.0168 0.0112 0.0321 0.0379 0.14637 0.12285
2D-CFD-IC (N) - - - - - - 0.08760 0.11385
2D-SW (N) 0.0044 0.083 0.0066 0.0024 0.00560 0.00657 0.00445 0.0021
3D-MHD (V) 0.36 0.1798 - - - - 0.3149 0.26845
3D-CFD (N) 0.37 1.0 0.299 - 0.262 - 0.09819 0.10573
Finetuning sets
1D-DR (N) 0.0014 0.006 - - - - 0.00158 0.00134
1D-BE (N) 0.029 0.37 - - - - 0.03961 0.07646
2D-CFD (N) 0.28 0.66 0.0312 0.0213 0.0176 0.0153 0.05318 0.05712
2D-GSDR (V) 0.1365 0.2252 - - - - 0.0086 0.01151
3D-CFD-Turb (N) 0.24 0.23 0.098 - - - 0.08297 0.07941
3D-TGC (V) 0.2429 0.6753 - - - - 0.16366 0.07153
Zero-shot on pretraining sets Zero-shot on finetuning sets
Datasets Models DPOT-FM Poseidon-FM MORPH-FM
S (30M) M (122M) T (21M) B (158M) Ti (7M) S (30M) M (126M) L* (480M)
FT sets (Full-shot)
1D-DR (N) - - - - 0.0008 0.00131 0.00173 0.00135
1D-BE (N) - - - - 0.0302 0.0584 0.04177 0.04123
3D-CFD-Turb (N) - - - - 0.10742 0.10205 0.08635 0.09170
3D-TGC (V) - - - - 0.05312 0.04136 0.03756 0.03246
Standalone surrogates.
In Table 3, we compare performance of MORPH standalone surrogates (MORPH-SS) on pretraining and finetuning datasets against existing baseline (FNO and UNets) and SOTA models (MPP and DPOT). The results for baselines and SOTA models are inherited from the respective benchmark sources (takamoto2022pdebench; ohana2024well) and articles (mccabe2024multiple; hao2024dpot). Our MORPH-SS-Ti and S models outperform baseline and SOTA models on standalone datasets and remain competitive across the rest. Practically, this makes MORPH a viable standalone surrogate when the target application focuses on a single PDE family.
Zero-shot performance.
Full-shot finetuning performance.
Full-shot fine-tuning results are presented in Table 5. On the 2D datasets, we compare the fine-tuning performance of MORPH-FM to DPOT-FM and Poseidon-FM, and find that our models consistently surpass these existing FMs. Together with Tables 3 and 5, we can quantify the adaptation-efficiency gains of MORPH-FM over MORPH-SS on the fine-tuning benchmarks: MORPH-FM outperforms MORPH-SS on 5 out of 6 fine-tuning datasets.
Parameter-efficient finetuning with LoRA.
We investigate fine-tuning the MORPH-FM-L model using low-rank adapters (LoRA). In our setup, only 77M of the 480M parameters comprising LoRA adapters on attention and MLP blocks, positional encodings and layer norms are fine-tuned, which is fewer trainable parameters than in the M model (126M). Notably, MORPH-FM-L matches or outperforms the M model across full-shot fine-tuning (Table 5). To our knowledge, this is one of the first two concurrent studies to apply LoRA to PDE foundation models (see also: holzschuh2025pde), and our results indicate that LoRA-based fine-tuning is a promising direction given the expanding PDE datasets and increasing model sizes. Future work will examine LoRA-based fine-tuning more extensively, including the effect of adapter rank across different datasets.
Autoregressive rollouts.
We have presented autoregressive rollouts results for standalone model and foundation model. Figs. 6 and 7 (Appendix C.3) show rollout results for MORPH-SS-Ti and MORPH-SS-S models on the Shallow Water Equations (SWE) dataset, using the t=0t{=}0 snapshot as input. Across the rollout horizon, MORPH-SS-S exhibits lower error than MORPH-SS-Ti, consistent with the NRMSE reported in Table 3. Both models produce stable multi-step forecasts, exhibiting limited error accumulation and no blow-up for the 10-step horizon. Fig. 3 presents autoregressive rollouts for MORPH-FM-S fine-tuned on the FNS-KF prediction task for $v_{y}$ ($v_{x}$ in Appendix C.3 Fig 9). It can be observed that MORPH-FM-S achieves better rollouts than MORPH-FM-Ti (Fig. 8 in Appendix C.3).
Figure 3: Finetuning MORPH-FM-S ($\sim$ 30M) for FNS-KF prediction task: 10-step autoregressive rollouts of $v_{y}$ with $t=0$ (initial frame) as input.
4.3 Scaling studies
Scaling w.r.t. finetuning dataset size.
Scaling w.r.t. finetuning model size (across FMs).
[…]
4.4 Ablation Studies
Model architecture.
Ablations Variant MHD3D DR2D CFD1D
Param. GFLOPs Loss Param. GFLOPs Loss Param. GFLOPs Loss
Conv Filters 0 (no conv) 8.29 0 0.18376 7.04 0 0.00787 7.14 0 0.17185
8 8.94 2.7556 0.16851 7.96 0.11376 0.00747 8.06 0.011 0.17348
Full 5.78 5.9056 0.01228 14.2 75.141 0.01208 4.67 0.6711 0.00291
Sparse 5.78 1.8832 0.01324 14.2 8.6064 0.00856 4.67 0.4372 0.00208
Axial 8.94 1.8832 0.00812 17.3 8.6064 0.00724 7.83 0.4372 0.00172
Pretraining corpus.
To evaluate the effect of the pretraining set and to disentangle the influence of the dataset from that of the model architecture, we pretrain MORPH-FM-S*(27M) on the same datasets as Poseidon. We then finetune both MORPH and Poseidon on 128 trajectories and test on 240 trajectories for 200 epochs across three downstream tasks: CE-RM, NS-PwC, and NS-SL. The results reported in Table 9 (Appendix C.2) demonstrate that MORPH-FM-S*(27M), with only one-fifth the number of parameters, performs competitively with or even outperforms Poseidon. These findings indicate that the MORPH architecture yields performance gains that are largely independent of the choice of pretraining corpus. More details about this study are provided in Appendix C.2, and full autoregressive rollouts are shown in Figs. 10, 12 and 11 (Appendix C.3).
5 Conclusion
We introduced MORPH, a modality-agnostic foundation model for PDEs that unifies heterogeneous spatiotemporal data across 1D–3D domains, scalar/vector fields, and mixed resolutions within a single convolutional vision transformer backbone. Our Unified Physics Tensor Format (UPTF-7) standardizes layout across PDE datasets while preserving physical semantics of the observations. By combining component-wise convolutions, inter-field cross-attention, and 4D axial attention, MORPH learns locality, cross-physics dependencies, and long-range spatiotemporal structure in a compute-aware way, and scales across multiple model variants. These design choices, together with parameter-efficient LoRA adapters for finetuning, enable effective transfer from diverse pretraining corpora to downstream tasks. We evaluated MORPH against its standalone version, task-specific baselines and recent PDE foundation models. Beyond prediction accuracy, MORPH highlights pathways toward flexible and scalable PDE foundation models for scientific machine learning.
Reproducibility statement
All datasets used for pretraining and fine-tuning are publicly available. We will release the full codebase and model checkpoints under an open-source license upon publication.
Acknowledgments
Appendices
Appendix A Data Details
A.1 Benchmark: PDEBench
A.1.1 1D Burgers’ equation
[…]
Category: Finetuning
A.1.2 1D Diffusion-reaction
PDE: $\partial_{t}u-\nu\,\partial_{xx}u-\rho\,u(1-u)=0$.
Domain: $x\in(0,1),\;t\in(0,1]$.
Category: Finetuning
A.1.3 2D Diffusion-reaction
Domain: $(x,y)\in(-1,1)^{2},\;t\in(0,5]$.
Constants: $k=5\times 10^{-3},D_{u}=1\times 10^{-3},D_{v}=5\times 10^{-3}$.
Numerical scheme Finite-volume in space and classical RK4 in time.
Category: Pretraining
A.1.4 Compressible Navier-Stokes (CNS) system
PDE:
$\partial_{t}\rho+\nabla\!\cdot(\rho\mathbf{v})$ $=0,$
BCs & ICs: Periodic BCs, Random ICs.
Numerical scheme: For advection/shock terms: HLLC + MUSCL (robust, shock-capturing, 2nd order). For viscous/diffusion terms: central differencing (smooth, consistent with Navier–Stokes viscosity). More details on the solver toro1994restoration; van1997towards.
Data specifications and category:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
A.1.5 2D inhomogeneous, incompressible Navier-Stokes
Domain: $\Omega\in[0,1]^{2}$.
Numerical scheme: Finite-difference simulations implemented in PhiFlow (holl2020phiflow).
Category: Pretraining
A.2 Benchmark: The Well
A.2.1 2D Gray-Scott diffusion reaction
Description: Two-species reaction–diffusion model whose scalar concentrations vary in space and time. It exhibits rich pattern formations and is used as a canonical test for nonlinear spatiotemporal dynamics and biological morphogenesis (gray1984autocatalytic).
Domain: $\Omega\in[-1,1]^{2}$.
BCs & ICs: Doubly periodic BCs
Constants: $\delta_{A}=2\times 10^{-5}$, $\delta_{B}=1\times 10^{-5}$.
Category: Finetuning.
A.2.2 3D Magnetohydrodynamics
Description: A multiphysics dataset coupling fluid dynamics and electromagnetism. MHD is a strongly nonlinear system in which coupled fluid–magnetic dynamics generate broadband turbulence, shocks and compressions, intermittent current sheets, and reconnection, yielding larger dynamics range of spatial and temporal scales. MHD modeling finds applications in understanding solar winds, interstellar medium, magnetospheres, fusion energy devices like tokamaks (burkhart2020catalogue).
PDE: Ideal MHD with isothermal EOS $p=c_{s}^{2}\rho$:
Domain: Periodic cube.
Numerical scheme: Third-order hybrid ENO shock-capturing scheme (cho2003compressible), isothermal EOS and forced turbulence. The high-res data were generated at $256^{3}$ and anti-aliased downsampled to $64^{3}$ for the MHD_64 split.
Category: Pretraining.
A.2.3 3D Turbulent cooling gravity
Domain: 3D volume.
Constants: Adiabatic index $\gamma=5/3$.
Numerical scheme: Density-Independent SPH (DISPH) in ASURA-FDPS; radiative processes via tabulated cooling/heating (hirashima20233d).
Category: Finetuning
A.3 Benchmark: PDEGYM
A.3.1 FNS-KF
PDE: Forced incompressible Navier–Stokes:
with time-independent forcing
$f(x,y)=0.1\,\sin\!\big(2\pi(x+y)\big),\qquad\partial_{t}f\equiv 0.$
Domain: $D=[0,1]^{2}$.
BCs & ICs: Periodic boundary conditions. Initial velocity is set exactly as in NS-PwC (herde2024poseidon) by prescribing a piecewise-constant vorticity $\omega_{0}$ on a $p\times p$ uniform partition (with $p=10$) and recovering a divergence-free velocity:
Data specifications: N=20, [^0] N=20{,}000 trajectories, $T=21$ time steps, $H\times W=128\times 128$, one vector field variables $(v)$. We assemble all the ‘.nc’ files in the data directory into a single combined file (herde2024poseidon). However, we use limited trajectories.
Category: Finetuning
A.4 Data Processing

### v4

@@ 4 Experiments
We select datasets from three publicly available PDE benchmarks, i.e., PDEBench (Takamoto et al., 2022), PDEgym (Herde et al., 2024) and The Well (Ohana et al., 2024). They cover a wide variety of PDEs; the latter, however, focuses on more complex multiphysics problems. Because MORPH accommodates data heterogeneity, we demonstrate versatility in our dataset selection for both pretraining and fine-tuning. The pretraining and finetuning datasets and their shapes are listed in Table 5 and 6 (Appendix A.4). Compared to existing PDE foundation models, our pretraining and fine-tuning sets are larger and more diverse. We defer more dataset-specific details to Appendix A.
Pretraining and finetuning settings.
Our model accommodates both fixed–order autoregressions, $AR(p)$, and time-varying orders, $AR(p_{1:T})$, where $p$ (or $p_{t}$) denotes the autoregressive order. In this paper, we train MORPH in a self-supervised $AR(1)$ setting where we learn a map $F_{\theta}:\mathcal{X}\to\mathcal{X}$ such that for all $t\geq 0$, $X_{t+1}\approx F_{\theta}(X_{t})$ by minimizing the mean-squared error. High–order fixed AR models (e.g., $AR(16)$ in McCabe et al. (2024) and $AR(10)$ in Hao et al. (2024)) often yield stronger long-horizon rollouts by reducing exposure bias via a longer receptive field. However, this design departs from the initial-value-problem (IVP) semantics that PDE foundation models aim to respect. For these reasons we adopt a one-step ($AR(1)$) formulation aligned with IVP semantics that preserves the compositional structure fundamental to PDE time evolution. Nevertheless, MORPH remains flexible and supports multi-order (fixed or varying) autoregression to improve rollout quality. For brevity, full pretraining and finetuning settings are documented in Appendix B.2, B.3.
4.1 Baselines
Models.
We compare MORPH against classical neural operators and recent PDE foundation models. As popular baselines, we use Fourier Neural Operator (FNO) (Li et al., 2020), a parameter-efficient spectral neural operator widely adopted for fluid surrogates, and UNet (Ronneberger et al., 2015; Rahman et al., 2022), a convolutional encoder–decoder with multiscale skip connections. Among foundation models, we include MPP (McCabe et al., 2024), which demonstrates gains from multi-physics pretraining; DPOT (Hao et al., 2024), a denoising operator transformer with Fourier attention; and Poseidon (Herde et al., 2024), a hierarchical multiscale operator transformer with shifted-window attention and lead-time-conditioned normalization for continuous-time querying.
Evaluation Metrics.
We follow the conventions of PDEBench (PB) and The Well (TW) when reporting errors on spatiotemporal PDE datasets. The Normalized Root Mean Squared Error (NRMSE) is the RMSE normalized by the $\ell_{2}$ norm of the ground truth, making it a scale-independent (dimensionless) measure and the default metric in PDEBench (Takamoto et al., 2022). In contrast, the Variance-Scaled RMSE (VRMSE) normalizes the RMSE by the target’s per-snapshot standard deviation (i.e., the square root of its variance), thereby comparing errors to the intrinsic variability of the field. The Well adopts VRMSE as a primary metric (Ohana et al., 2024).
Comparison Caveats.
4.2 Main results
Our primary experiments are organized in Tables 1,2,3. In these tables, we compare baseline and SOTA models using NRMSE and VRMSE as evaluation metrics.
Datasets Models FNO UNet MPP-$S$ DPOT-$S$ MORPH-$S$
PB/TW PB/TW Ti (7M) S (30M) Ti (7M) S (30M) Ti (7M) S (30M)
Pretraining sets
1D-CFD (N) 0.095 0.36 - - - - 0.06102 0.05719
2D-DR (N) 0.12 0.84 0.0168 0.0112 0.0321 0.0379 0.14637 0.12285
2D-CFD-IC (N) - - - - - - 0.08760 0.11385
2D-SW (N) 0.0044 0.083 0.0066 0.0024 0.00560 0.00657 0.00445 0.0021
3D-MHD (V) 0.36 0.1798 - - - - 0.3149 0.26845
3D-CFD (N) 0.37 1.0 0.299 - 0.262 - 0.09819 0.10573
Finetuning sets
1D-DR (N) 0.0014 0.006 - - - - 0.00158 0.00134
1D-BE (N) 0.029 0.37 - - - - 0.03961 0.07646
2D-CFD (N) 0.28 0.66 0.0312 0.0213 0.0176 0.0153 0.05318 0.05712
2D-GSDR (V) 0.1365 0.2252 - - - - 0.0086 0.01151
3D-CFD-Turb (N) 0.24 0.23 0.098 - - - 0.08297 0.07941
3D-TGC (V) 0.2429 0.6753 - - - - 0.16366 0.07153
Zero-shot on pretraining sets Zero-shot on finetuning sets
Datasets Models DPOT-FM Poseidon-FM MORPH-FM
S (30M) M (122M) T (21M) B (158M) Ti (7M) S (30M) M (126M) L* (480M)
FT sets (Full-shot)
1D-DR (N) - - - - 0.0008 0.00131 0.00173 0.00135
1D-BE (N) - - - - 0.0302 0.0584 0.04177 0.04123
3D-CFD-Turb (N) - - - - 0.10742 0.10205 0.08635 0.09170
3D-TGC (V) - - - - 0.05312 0.04136 0.03756 0.03246
Standalone surrogates.
In Table 1, we compare performance of MORPH standalone surrogates (MORPH-SS) on pretraining and finetuning datasets against existing baseline (FNO and UNets) and SOTA models (MPP and DPOT). The results for baselines and SOTA models are inherited from the respective benchmark sources (Takamoto et al., 2022; Ohana et al., 2024) and articles (McCabe et al., 2024; Hao et al., 2024). Our MORPH-SS-Ti and S models outperform baseline and SOTA models on standalone datasets and remain competitive across the rest. Practically, this makes MORPH a viable standalone surrogate when the target application focuses on a single PDE family.
Zero-shot performance.
Full-shot finetuning performance.
Full-shot fine-tuning results are presented in Table 3. On the 2D datasets, we compare the fine-tuning performance of MORPH-FM to DPOT-FM and Poseidon-FM, and find that our models consistently surpass these existing FMs. Together with Tables 1 and 3, we can quantify the adaptation-efficiency gains of MORPH-FM over MORPH-SS on the fine-tuning benchmarks: MORPH-FM outperforms MORPH-SS on 5 out of 6 fine-tuning datasets.
Parameter-efficient finetuning with LoRA.
We investigate fine-tuning the MORPH-FM-L model using low-rank adapters (LoRA). In our setup, only 77M of the 480M parameters comprising LoRA adapters on attention and MLP blocks, positional encodings and layer norms are fine-tuned, which is fewer trainable parameters than in the M model (126M). Notably, MORPH-FM-L matches or outperforms the M model across full-shot fine-tuning (Table 3). To our knowledge, this is one of the first two concurrent studies to apply LoRA to PDE foundation models (see also: (Holzschuh et al., 2025)), and our results indicate that LoRA-based fine-tuning is a promising direction given the expanding PDE datasets and increasing model sizes. Future work will examine LoRA-based fine-tuning more extensively, including the effect of adapter rank across different datasets.
Autoregressive rollouts.
We have presented autoregressive rollouts results for standalone model and foundation model. Figs. 5 and 6 (Appendix C.3) show rollout results for MORPH-SS-Ti and MORPH-SS-S models on the Shallow Water Equations (SWE) dataset, using the t=0t{=}0 snapshot as input. Across the rollout horizon, MORPH-SS-S exhibits lower error than MORPH-SS-Ti, consistent with the NRMSE reported in Table 1. Both models produce stable multi-step forecasts, exhibiting limited error accumulation and no blow-up for the 10-step horizon. Fig 8 (Appendix C.3) presents autoregressive rollouts for MORPH-FM-S fine-tuned on the FNS-KF prediction task. It can be observed that MORPH-FM-S achieves better rollouts than MORPH-FM-Ti (Fig. 7 in Appendix C.3), which is attributed to its larger capacity (i.e., more parameters). Further rollout studies are discussed in following sections.
4.3 Scaling studies
Scaling w.r.t. finetuning dataset size.
Scaling w.r.t. finetuning model size (across FMs).
[…]
4.4 Ablation Studies
Ablations Variant MHD3D DR2D CFD1D
Param. GFLOPs Loss Param. GFLOPs Loss Param. GFLOPs Loss
Conv Filters 0 (no conv) 8.29 0 0.18376 7.04 0 0.00787 7.14 0 0.17185
8 8.94 2.7556 0.16851 7.96 0.11376 0.00747 8.06 0.011 0.17348
Full 5.78 5.9056 0.01228 14.2 75.141 0.01208 4.67 0.6711 0.00291
Sparse 5.78 1.8832 0.01324 14.2 8.6064 0.00856 4.67 0.4372 0.00208
Axial 8.94 1.8832 0.00812 17.3 8.6064 0.00724 7.83 0.4372 0.00172
Model architecture.
Pretraining corpus.
To evaluate the effect of the pretraining set and to disentangle the influence of the dataset from that of the model architecture, we pretrain MORPH-FM-S*(27M) on the same datasets as Poseidon. We then finetune both MORPH and Poseidon on 128 trajectories and test on 240 trajectories for 200 epochs across three downstream tasks: CE-RM, NS-PwC, and NS-SL. The results reported in Table 9 (Appendix C.2) demonstrate that MORPH-FM-S*(27M), with only one-fifth the number of parameters, performs competitively with or even outperforms Poseidon. Figs. 9, 10 and 11 (Appendix C.3) show show full autoregressive rollouts on representative test trajectories for NS-SL, NS-PwC, and CE-RM. These findings indicate that the MORPH architecture yields performance gains that are largely independent of the choice of pretraining corpus. Additional details about this study are provided in Appendix C.2.
5 Conclusion
We introduced MORPH, a modality-agnostic foundation model for PDEs that unifies heterogeneous spatiotemporal data across 1D–3D domains, scalar/vector fields, and mixed resolutions within a single convolutional vision transformer backbone. Our Unified Physics Tensor Format (UPTF-7) standardizes layout across PDE datasets while preserving physical semantics of the observations. Model design choices, together with parameter-efficient adapters for finetuning, enable effective transfer from diverse pretraining corpora to downstream tasks. We evaluated MORPH against its standalone version, task-specific baselines and recent PDE foundation models. Beyond prediction accuracy, MORPH highlights pathways toward flexible and scalable PDE foundation models for scientific machine learning.
Impact Statement
MORPH enables modality-agnostic PDE surrogate modeling across heterogeneous 1D–3D spatiotemporal data, reducing reliance on expensive simulation, experiments and task-specific model redesign. It can accelerate scientific workflows by learning from limited observations and transferring to new tasks with parameter-efficient adaptations. In particular, MORPH can support both forward and inverse problems within scientific machine learning pipelines. Potential risks include dual-use (speeding up sensitive engineering applications) and harm from over-trusting inaccurate predictions in safety-critical settings. MORPH should be applied with careful domain validation and uncertainty-aware evaluation.
Acknowledgments
Appendices
Appendix A Data Details
A.1 Benchmark: PDEBench
A.1.1 1D Burgers’ equation
[…]
Category: Finetuning
A.1.2 1D Diffusion-reaction
PDE: $\partial_{t}u-\nu\,\partial_{xx}u-\rho\,u(1-u)=0$.
Domain: $x\in(0,1),\;t\in(0,1]$.
Category: Finetuning
A.1.3 2D Diffusion-reaction
Domain: $(x,y)\in(-1,1)^{2},\;t\in(0,5]$.
Constants: $k=5\times 10^{-3},D_{u}=1\times 10^{-3},D_{v}=5\times 10^{-3}$.
Numerical scheme Finite-volume in space and classical RK4 in time.
Category: Pretraining
A.1.4 Compressible Navier-Stokes (CNS) system
PDE:
$\partial_{t}\rho+\nabla\!\cdot(\rho\mathbf{v})$ $=0,$
BCs & ICs: Periodic BCs, Random ICs.
Numerical scheme: For advection/shock terms: HLLC + MUSCL (robust, shock-capturing, 2nd order). For viscous/diffusion terms: central differencing (smooth, consistent with Navier–Stokes viscosity). More details on the solver Toro et al. (1994); Van Leer (1997).
Data specifications and category:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
A.1.5 2D inhomogeneous, incompressible Navier-Stokes
Domain: $\Omega\in[0,1]^{2}$.
Numerical scheme: Finite-difference simulations implemented in PhiFlow (Holl et al., 2020).
Category: Pretraining
A.2 Benchmark: The Well
A.2.1 2D Gray-Scott diffusion reaction
Description: Two-species reaction–diffusion model whose scalar concentrations vary in space and time. It exhibits rich pattern formations and is used as a canonical test for nonlinear spatiotemporal dynamics and biological morphogenesis (Gray and Scott, 1984).
Domain: $\Omega\in[-1,1]^{2}$.
BCs & ICs: Doubly periodic BCs
Constants: $\delta_{A}=2\times 10^{-5}$, $\delta_{B}=1\times 10^{-5}$.
Category: Finetuning.
A.2.2 3D Magnetohydrodynamics
Description: A multiphysics dataset coupling fluid dynamics and electromagnetism. MHD is a strongly nonlinear system in which coupled fluid–magnetic dynamics generate broadband turbulence, shocks and compressions, intermittent current sheets, and reconnection, yielding larger dynamics range of spatial and temporal scales. MHD modeling finds applications in understanding solar winds, interstellar medium, magnetospheres, fusion energy devices like tokamaks (Burkhart et al., 2020).
PDE: Ideal MHD with isothermal EOS $p=c_{s}^{2}\rho$:
Domain: Periodic cube.
Numerical scheme: Third-order hybrid ENO shock-capturing scheme (Cho and Lazarian, 2003), isothermal EOS and forced turbulence. The high-res data were generated at $256^{3}$ and anti-aliased downsampled to $64^{3}$ for the MHD_64 split.
Category: Pretraining.
A.2.3 3D Turbulent cooling gravity
Domain: 3D volume.
Constants: Adiabatic index $\gamma=5/3$.
Numerical scheme: Density-Independent SPH (DISPH) in ASURA-FDPS; radiative processes via tabulated cooling/heating (Hirashima et al., 2023).
Category: Finetuning
A.3 Benchmark: PDEGYM
A.3.1 FNS-KF
PDE: Forced incompressible Navier–Stokes:
with time-independent forcing
$f(x,y)=0.1\,\sin\!\big(2\pi(x+y)\big),\qquad\partial_{t}f\equiv 0.$
Domain: $D=[0,1]^{2}$.
BCs & ICs: Periodic boundary conditions. Initial velocity is set exactly as in NS-PwC (Herde et al., 2024) by prescribing a piecewise-constant vorticity $\omega_{0}$ on a $p\times p$ uniform partition (with $p=10$) and recovering a divergence-free velocity:
Data specifications: N=20, [^0] N=20{,}000 trajectories, $T=21$ time steps, $H\times W=128\times 128$, one vector field variables $(v)$. We assemble all the ‘.nc’ files in the data directory into a single combined file (Herde et al., 2024). However, we use limited trajectories.
Category: Finetuning
A.4 Pretraining and finetuning sets
The pretraining and fine-tuning datasets, along with their field compositions and tensor shapes, are summarized in Tables 5 and 6. Collectively, this suite spans 1D–3D modalities, resolutions from 1D grids of length 1024 to 2D grids up to $512{\times}512$ and 3D volumes up to $128^{3}$, and includes both scalar and vector fields across single- and multi-physics systems.
@@ A.4 Pretraining and finetuning sets
A.5 Data Processing